Harris Company manufactures and sells a single product. A partially completed schedule of the company’s total costs and costs per unit over the relevant range of 50,000 to 90,000 units is given below: Required: 1. Complete the schedule of the company’s total costs and costs per unit as given in the relevant tab below. 2. Assume that the company produces and sells 80,000 units during the year at a selling price of $9.77 per unit. Prepare a contribution format income statement for the year.

Complete the schedule of the company’s total costs and costs per unit as given in the relevant tab below. (Round the per unit variable cost and fixed cost to 2 decimal places.)

Units Produced and Sold
50,000 70,000 90,000
Total costs:
Variable costs $160,000
Fixed costs 390,000
Total costs $550,000 $0 $0
Cost per unit:
Variable cost
Fixed cost
Total cost per unit $0.00 $0.00 $0.00

Assume that the company produces and sells 80,000 units during the year at a selling price of $9.77 per unit. Prepare a contribution format income statement for the year.
Answer #1


Harris Company Units Produced and Sold 50000 70000 90000 160000 390000 550000 224000 390000 614000 288000 390000 678000 Total

А 1 Harris Company 2 3 Units Produced and sold 70000 90000 50000 160000 390000 =SUM(B6:37) =C4*010 390000 =SUM(C6:07) =D4*D10